# Scalable Bayesian Learning with posteriors

### v1

@@ 7 Conclusion and outlook
We have presented a new software package posteriors for scalable and extensible uncertainty quantification with PyTorch [62] models. Through numerical experiments, we have demonstrated how posteriors can be used to achieve a range of benefits over optimization-based approaches and composing with PyTorch toolkits. In particular, we highlighted that the cold posterior effect mostly occurs in Gaussian approximations to the Bayesian posterior as well as demonstrating the benefits of Bayesian learning in the very practical and large-scale setting of free-form text generation with large language models.
We are excited by the future prospects for posteriors. Its functional API makes it easy to maintain and extend; we already have a long list of methods and approaches to add. posteriors is fully open source, and we are keen to foster a community for feedback for feature requests and encourage contributions. The framework from Section 4 allows for simultaneous research into optimization and sampling methods. Combining that with posteriors’ broad range of utility functions (that, to our knowledge, are not available elsewhere in the PyTorch ecosystem), means that implementation and research into second-order techniques [56, 54] and improved discretization methods [44] represent accessible avenues for future work in both optimization and SGMCMC.
Acknowledgments and Disclosure of Funding
We thank Patrick Coles, Antonio Martinez and Samuel Power for helpful comments on the manuscript. We also thank Dmitri Saberi for support in the development of posteriors.
@@ A.1 Laplace approximation
As discussed, the Laplace approximation [50, 15] forms a Gaussian approximation to the posterior distribution $p(\theta\mid y_{1:N})\approx\operatorname{\mathbf{N}}(\theta\mid\theta_{\text{MAP}},\hat{\Sigma})$ based on the second order Taylor expansion around the posterior mode $\theta_{\text{MAP}}=\operatorname*{arg\,max}_{\theta}p(\theta\mid y_{1:N})$:
@@
The Taylor expansion implies the use of the negative inverse Hessian $-\nabla^{2}_{\theta}\log p(\theta_{\text{MAP}}\mid y_{1:N})^{-1}$ as the covariance $\hat{\Sigma}$, however this is not guaranteed to be positive semidefinite and is therefore invalid as a covariance matrix. Instead, we can draw from the Bernstein-von Mises theorem [29]:
@@
$\operatorname{\mathrm{F}}(\theta)=\operatorname{\mathbb{E}}_{p(y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid y)]=\operatorname{\mathbb{E}}_{p(y\mid\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}].$ (6)
In the supervised learning setting with inputs $x$, the Fisher information matrix becomes $\operatorname{\mathrm{F}}(\theta)=\operatorname{\mathbb{E}}_{p(x,y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid x,y)]$ which is intractable as we do not have access to the distribution over inputs $p(x)$. Two popular approximations use the training data $\hat{p}(x,y)=N^{-1}\sum_{i=1}^{N}\delta(x,y\mid x^{i},y^{i})$. The empirical Fisher:
$\operatorname{\mathrm{F}}_{\text{E}}(\theta)=\operatorname{\mathbb{E}}_{\hat{p}(x,y)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}],$
@@
$\operatorname{\mathrm{F}}_{\text{C}}(\theta)=\operatorname{\mathbb{E}}_{\hat{p}(x)p(y\mid x,\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}],$ (7)
with the conditional Fisher generally recommended [42].
The integral with respect to $p(y\mid x,\theta)$ in (7) may initially seem difficult to compute. Fortunately, things are simplified in the common machine learning setting where the likelihood has the form $p(y\mid x,\theta)=p(y\mid f_{\theta}(x))$ where $p(y\mid z)$ is the probability density function for an exponential family distribution with natural parameter $z$. That is
@@
for some $T(y)$. This setting includes the common cases of classification (softmax with cross-entropy loss) and regression (mean squared error loss). Then we have an exact equivalence between $F_{C}(\theta)$ and the Generalized Gauss-Newton matrix (GGN) [54] matrix:
@@
Laplace approximation training consists of first determining the posterior mode $\theta_{\text{MAP}}$ through gradient descent. This is followed by a single additional epoch to calculate the approximate covariance. It is, thus, cheap to compute in the large data regime. However, in the large parameter regime, storing the covariance matrix with $\operatorname{\mathcal{O}}(d^{2})$ elements becomes a memory bottleneck as well as the required inverting and sampling operations, which typically have runtime $\operatorname{\mathcal{O}}(d^{3})$. As such, approximations are used, such as diagonal [50], KFAC [68] and low-rank [15] with reduced $\operatorname{\mathcal{O}}(d)$ memory. The main caveat for Laplace approximations is the accuracy of a Gaussian approximation under the extreme non-linearities in a neural network posterior [35], which applies even further to reduced memory approximations.
@@ A.2 Variational inference
Variational inference seeks a parameterized distribution $q_{\phi}(\theta)$ that minimizes some distance from the target posterior distribution $p(\theta\mid y_{1:N})$. Most commonly [7] we assume a Gaussian variational distribution $q_{\phi}(\theta)=\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$ with $\phi=(\mu,\Sigma)$ and the distance to be minimized is the KL divergence:
$0\leq\text{KL}[q_{\phi}(\theta)\mid\mid p(\theta\mid y_{1:N})]=\log p(y_{1:N})+\underbrace{\operatorname{\mathbb{E}}_{q_{\phi}(\theta)}\left[\log\frac{q_{\phi}(\theta)}{p(\theta)}-\log p(y_{1:N}\mid\theta)\right]}_{\text{NELBO}(\phi)}.$ (8)
Thus, the problem becomes one of optimization to minimize the negative evidence lower bound objective $\text{NELBO}(\phi)$, which can be achieved with stochastic gradient descent alongside the reparameterization trick [67] to obtain unbiased gradients of the expectation.
@@
Notably, the $\text{NELBO}(\phi)$ objective can be estimated stochastically by sampling from $q_{\phi}(\theta)$ and even minibatching the data $y_{1:N}$. Therefore, training is scalable in the large data regime. The caveat regarding the crudeness of a Gaussian approximation to the posterior also applies here, except that variational inference minimizes a formal divergence from the posterior rather than relying on Bernstein-von Mises asymptotics. However, in the large parameter regime, the $\text{NELBO}(\phi)$ objective requires high dimensional integration at every optimization step, which can suffer from high variance even with the available tricks [7].
@@ A.3 Stochastic gradient Markov chain Monte Carlo
A completely different paradigm to posterior approximation is to construct a Monte Carlo approximation $p(\theta\mid y_{1:N})\approx N^{-1}\sum_{k=1}^{K}\delta(\theta\mid\theta_{k})$. This is most commonly achieved by evolving a stochastic differential equation (SDE) and collecting samples along the trajectory. Following [48], an SDE has $\pi(z)$ as a stationary distribution (and therefore provides an asymptotically unbiased Monte Carlo approximation) if and only if it is of the form
@@
with symmetric positive-semidefinite $\operatorname{\mathrm{D}}(z)$ and skew symmetric $\operatorname{\mathrm{Q}}(z)=-\operatorname{\mathrm{Q}}(z)^{\top}$. Here $z=(\theta,\omega)$ where $\omega$ are any auxiliary parameters such as momenta and thermostat [48, 18]. The distribution $\pi(z)$ is thus an extended target with $p(\theta\mid y_{1:N})$ marginal in $\theta$. The matrices $\operatorname{\mathrm{D}}(z)$ and $\operatorname{\mathrm{Q}}(z)$ are often chosen as independent of $z$, meaning the matrix divergence term vanishes.
The exact simulation of such a non-linear SDE is intractable, and as such, approximate discretization schemes such as the Euler-Maruyama [76] method are employed. Traditional full-batch Markov chain Monte Carlo (MCMC) schemes [4] use a Metropolis-Hastings step to ensure convergence of the Monte Carlo approximation is preserved through discretization. In the minibatch setting, the Metropolis-Hastings ratio cannot be estimated easily. Instead, we can use an unbiased stochastic gradient $\widehat{\nabla\log\pi}(z)\sim\operatorname{\mathbf{N}}(\cdot\mid\nabla\log\pi(z),\mathrm{B}(z))$, then as noted in [76] the stochasticity arising from minibatching, $\mathrm{B}(z)$, decreases like $\operatorname{\mathcal{O}}(\epsilon_{t}^{2})$ for learning rate $\epsilon_{t}$. A Robbins-Monro [69] learning rate schedule $\sum\epsilon_{t}^{2}<\sum\epsilon_{t}=\infty$ ensures an asymptotically unbiased Monte Carlo approximation [76]. Additional enhancements can be made if an approximation to $\mathrm{B}(z)$ is available [48].
@@
The stochastic gradient formulation of MCMC permits easy application to the minibatch and large data regime. However, the decreasing learning rate requirement results in diffusive behaviour [3, 45, 5] where the discretized SDE puts excessive weight on the noise term $dw$ over the informative gradient term. This is problematic since exploration is very much desired for MCMC. Often, a constant learning rate is used in practice [44] with an asymptotic bias incurred. In the large parameter regime, the formulation of a Monte Carlo approximation becomes a memory bottleneck. Typical lower dimensional MCMC procedures take many thousands of samples, but due to memory constraints, only a significantly smaller sample size is available for large neural network parameters, resulting in a crude approximation. A notable further caveat is that, in contrast to the Gaussian approximations, the resulting Monte Carlo approximation does not provide access to a density with pointwise evaluations and cannot be used as a prior for subsequent Bayesian updates in an online regime (1).
@@ A.4 Notable mentions
Perhaps the most common route to uncertainty quantification in neural networks is dropout [25, 26], where model parameters or neurons are randomly removed from each forward call. This approach can be effective and has a Bayesian interpretation in the posterior probability space of inclusion/exclusion, but this is naturally more restricted than the posterior approximations above.
Stochastic weight averaging (SWA) [34] and SWA-Gaussian (SWAG) [51] are similar in spirit to SGMCMC as they use samples from a single trajectory to generate a point or Gaussian approximation to the posterior respectively. Another option for formulating a Gaussian distribution over parameters is to consider each minibatch as a small-scale Bayesian update; local linearization of this update is the paradigm of the extended Kalman filter [21], which also has natural gradient descent [60] and variational inference interpretations [82]. These methods are easy to implement but again suffer from the caveats of Gaussian posterior approximations above.
@@ Appendix B SGMCMC algorithms
Recall the generic SDE [48] from Section 4 for targeting the tempered $\pi_{\operatorname{\mathcal{T}}}(z)\propto\pi(z)^{\frac{1}{N\operatorname{\mathcal{T}}}}$ where we set $\operatorname{\mathcal{T}}=N^{-1}$ to target the Bayesian posterior
@@
The matrix divergence $\nabla\cdot\mathrm{M}(z)\in\operatorname{\mathbf{R}}^{d}$ is defined [48] as
@@
We now cover the most popular SGMCMC algorithms and those implemented in posteriors (all of which are also found in [48]).
@@ B.1 SGLD
The simplest SGMCMC algorithm is stochastic gradient Langevin dynamics [76], which contains no auxiliary variables (i.e. $z=\theta$) and is represented by $\operatorname{\mathrm{D}}=\mathbb{I}$ and $\operatorname{\mathrm{Q}}=0$. This gives continuous-time SGLD:
@@ B.2 SGHMC
Stochastic gradient Hamiltonian Monte Carlo [12] extends $z=(\theta,m)$ to include momenta $m$ with Gaussian target distribution
@@ B.3 SGNHT
The stochastic gradient Nosé-Hoover thermostat (SGNHT) algorithm [18] uses ideas from thermodynamics [44] to introduce an additional scalar variable $\xi$ that automates the selection of the $\alpha$ friction parameter in SGHMC. We have $z=(\theta,m,\xi)$ targeting  [^2] Note that the SGNHT target distribution for $\xi$ is erroneously claimed as $\frac{1}{2d}(\xi-\alpha)^{2}$ in [48] and [58]

### v2

@@ 7 Conclusion and outlook
We have presented a new software package posteriors for scalable and extensible uncertainty quantification with PyTorch (Paszke et al., 2019) models. Through numerical experiments, we have demonstrated how posteriors can be used to achieve a range of benefits over optimization-based approaches and composing with PyTorch toolkits. In particular, we highlighted that the cold posterior effect mostly occurs in Gaussian approximations to the Bayesian posterior as well as demonstrating the benefits of Bayesian learning in the very practical and large-scale setting of free-form text generation with large language models.
We are excited by the future prospects for posteriors. Its functional API makes it easy to maintain and extend; we already have a long list of methods and approaches to add. posteriors is fully open source, and we are keen to foster a community for feedback for feature requests and encourage contributions. The framework from Section 4 allows for simultaneous research into optimization and sampling methods. Combining that with posteriors’ broad range of utility functions (that, to our knowledge, are not available elsewhere in the PyTorch ecosystem), means that implementation and research into second-order techniques (Martens et al., 2010; Martens, 2020) and improved discretization methods (Leimkuhler et al., 2018) represent accessible avenues for future work in both optimization and SGMCMC.
@@ A.1 Laplace approximation
As discussed, the Laplace approximation [MacKay, 1992, Daxberger et al., 2021] forms a Gaussian approximation to the posterior distribution $p(\theta\mid y_{1:N})\approx\operatorname{\mathbf{N}}(\theta\mid\theta_{\text{MAP}},\hat{\Sigma})$ based on the second order Taylor expansion around the posterior mode $\theta_{\text{MAP}}=\operatorname*{arg\,max}_{\theta}p(\theta\mid y_{1:N})$:
@@
The Taylor expansion implies the use of the negative inverse Hessian $-\nabla^{2}_{\theta}\log p(\theta_{\text{MAP}}\mid y_{1:N})^{-1}$ as the covariance $\hat{\Sigma}$, however this is not guaranteed to be positive semidefinite and is therefore invalid as a covariance matrix. Instead, we can draw from the Bernstein-von Mises theorem [Hartigan and Hartigan, 1983]:
@@
$\operatorname{\mathrm{F}}(\theta)=\mathbb{E}_{p(y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid y)]=\mathbb{E}_{p(y\mid\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}].$ (6)
In the supervised learning setting with inputs $x$, the Fisher information matrix becomes $\operatorname{\mathrm{F}}(\theta)=\mathbb{E}_{p(x,y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid x,y)]$ which is intractable as we do not have access to the distribution over inputs $p(x)$. Two popular approximations use the training data $\hat{p}(x,y)=N^{-1}\sum_{i=1}^{N}\delta(x,y\mid x^{i},y^{i})$. The empirical Fisher:
$\operatorname{\mathrm{F}}_{\text{E}}(\theta)=\mathbb{E}_{\hat{p}(x,y)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}],$
@@
$\operatorname{\mathrm{F}}_{\text{C}}(\theta)=\mathbb{E}_{\hat{p}(x)p(y\mid x,\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}],$ (7)
with the conditional Fisher generally recommended [Kunstner et al., 2019].
The integral with respect to $p(y\mid x,\theta)$ in equation 7 may initially seem difficult to compute. Fortunately, things are simplified in the common machine learning setting where the likelihood has the form $p(y\mid x,\theta)=p(y\mid f_{\theta}(x))$ where $p(y\mid z)$ is the probability density function for an exponential family distribution with natural parameter $z$. That is
@@
for some $T(y)$. This setting includes the common cases of classification (softmax with cross-entropy loss) and regression (mean squared error loss). Then we have an exact equivalence between $F_{C}(\theta)$ and the Generalized Gauss-Newton matrix (GGN, Martens [2020]) matrix:
@@
Laplace approximation training consists of first determining the posterior mode $\theta_{\text{MAP}}$ through gradient descent. This is followed by a single additional epoch to calculate the approximate covariance. It is, thus, cheap to compute in the large data regime. However, in the large parameter regime, storing the covariance matrix with $\operatorname{\mathcal{O}}(d^{2})$ elements becomes a memory bottleneck as well as the required inverting and sampling operations, which typically have runtime $\operatorname{\mathcal{O}}(d^{3})$. As such, approximations are used, such as diagonal [MacKay, 1992], K-FAC [Ritter et al., 2018] and low-rank [Daxberger et al., 2021] with reduced $\operatorname{\mathcal{O}}(d)$ memory. The main caveat for Laplace approximations is the accuracy of a Gaussian approximation under the extreme non-linearities in a neural network posterior [Izmailov et al., 2021], which applies even further to reduced memory approximations.
@@ A.2 Variational inference
Variational inference seeks a parameterized distribution $q_{\phi}(\theta)$ that minimizes some distance from the target posterior distribution $p(\theta\mid y_{1:N})$. Most commonly [Blei et al., 2017] we assume a Gaussian variational distribution $q_{\phi}(\theta)=\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$ with $\phi=(\mu,\Sigma)$ and the distance to be minimized is the KL divergence:
$0\leq\text{KL}[q_{\phi}(\theta)\mid\mid p(\theta\mid y_{1:N})]=\log p(y_{1:N})+\underbrace{\mathbb{E}_{q_{\phi}(\theta)}\left[\log\frac{q_{\phi}(\theta)}{p(\theta)}-\log p(y_{1:N}\mid\theta)\right]}_{\text{NELBO}(\phi)}.$ (8)
Thus, the problem becomes one of optimization to minimize the negative evidence lower bound objective $\text{NELBO}(\phi)$, which can be achieved with stochastic gradient descent alongside the reparameterization trick [Rezende et al., 2014] to obtain unbiased gradients of the expectation.
@@
Notably, the $\text{NELBO}(\phi)$ objective can be estimated stochastically by sampling from $q_{\phi}(\theta)$ and even minibatching the data $y_{1:N}$. Therefore, training is scalable in the large data regime. The caveat regarding the crudeness of a Gaussian approximation to the posterior also applies here, except that variational inference minimizes a formal divergence from the posterior rather than relying on Bernstein-von Mises asymptotics. However, in the large parameter regime, the $\text{NELBO}(\phi)$ objective requires high dimensional integration at every optimization step, which can suffer from high variance even with the available tricks [Blei et al., 2017].
@@ A.3 Stochastic gradient Markov chain Monte Carlo
A completely different paradigm to posterior approximation is to construct a Monte Carlo approximation $p(\theta\mid y_{1:N})\approx N^{-1}\sum_{k=1}^{K}\delta(\theta\mid\theta_{k})$. This is most commonly achieved by evolving a stochastic differential equation (SDE) and collecting samples along the trajectory. Following Ma et al. [2015], an SDE has $\pi(z)$ as a stationary distribution (and therefore provides an asymptotically unbiased Monte Carlo approximation) if and only if it is of the form
@@
with symmetric positive-semidefinite $\operatorname{\mathrm{D}}(z)$ and skew symmetric $\operatorname{\mathrm{Q}}(z)=-\operatorname{\mathrm{Q}}(z)^{\top}$. Here $z=(\theta,\omega)$ where $\omega$ are any auxiliary parameters such as momenta and thermostat [Ma et al., 2015, Ding et al., 2014]. The distribution $\pi(z)$ is thus an extended target with $p(\theta\mid y_{1:N})$ marginal in $\theta$. The matrices $\operatorname{\mathrm{D}}(z)$ and $\operatorname{\mathrm{Q}}(z)$ are often chosen as independent of $z$, meaning the matrix divergence term vanishes.
The exact simulation of such a non-linear SDE is intractable, and as such, approximate discretization schemes such as the Euler-Maruyama [Welling and Teh, 2011] method are employed. Traditional full-batch Markov chain Monte Carlo (MCMC) schemes [Andrieu et al., 2003] use a Metropolis-Hastings step to ensure convergence of the Monte Carlo approximation is preserved through discretization. In the minibatch setting, the Metropolis-Hastings ratio cannot be estimated easily. Instead, we can use an unbiased stochastic gradient $\widehat{\nabla\log\pi}(z)\sim\operatorname{\mathbf{N}}(\cdot\mid\nabla\log\pi(z),\mathrm{B}(z))$, then as noted in Welling and Teh [2011] the stochasticity arising from minibatching, $\mathrm{B}(z)$, decreases like $\operatorname{\mathcal{O}}(\epsilon_{t}^{2})$ for learning rate $\epsilon_{t}$. A Robbins-Monro [Robbins and Monro, 1951] learning rate schedule $\sum\epsilon_{t}^{2}<\sum\epsilon_{t}=\infty$ ensures an asymptotically unbiased Monte Carlo approximation [Welling and Teh, 2011]. Additional enhancements can be made if an approximation to $\mathrm{B}(z)$ is available [Ma et al., 2015].
@@
The stochastic gradient formulation of MCMC permits easy application to the minibatch and large data regime. However, the decreasing learning rate requirement results in diffusive behaviour [Alexos et al., 2022, Li et al., 2016, Bieringer et al., 2023] where the discretized SDE puts excessive weight on the noise term $dw$ over the informative gradient term. This is problematic since exploration is very much desired for MCMC. Often, a constant learning rate is used in practice [Leimkuhler et al., 2018] with an asymptotic bias incurred. In the large parameter regime, the formulation of a Monte Carlo approximation becomes a memory bottleneck. Typical lower dimensional MCMC procedures take many thousands of samples, but due to memory constraints, only a significantly smaller sample size is available for large neural network parameters, resulting in a crude approximation. A notable further caveat is that, in contrast to the Gaussian approximations, the resulting Monte Carlo approximation does not provide access to a density with pointwise evaluations and cannot be used as a prior for subsequent Bayesian updates in an online regime equation 1.
@@ A.4 Notable mentions
Perhaps the most common route to uncertainty quantification in neural networks is dropout [Gal and Ghahramani, 2016, Gal et al., 2017], where model parameters or neurons are randomly removed from each forward call. This approach can be effective and has a Bayesian interpretation in the posterior probability space of inclusion/exclusion, but this is naturally more restricted than the posterior approximations above.
Stochastic weight averaging (SWA, Izmailov et al. [2018]) and SWA-Gaussian (SWAG, Maddox et al. [2019]) are similar in spirit to SGMCMC as they use samples from a single trajectory to generate a point or Gaussian approximation to the posterior respectively. Another option for formulating a Gaussian distribution over parameters is to consider each minibatch as a small-scale Bayesian update; local linearization of this update is the paradigm of the extended Kalman filter [Duran-Martin et al., 2022], which also has natural gradient descent [Ollivier, 2018] and variational inference interpretations [Zhang et al., 2018]. These methods are easy to implement but again suffer from the caveats of Gaussian posterior approximations above.
@@ Appendix B SGMCMC algorithms
Recall the generic SDE Ma et al. [2015] from Section 4 for targeting the tempered $\pi_{\operatorname{\mathcal{T}}}(z)\propto\pi(z)^{\frac{1}{N\operatorname{\mathcal{T}}}}$ where we set $\operatorname{\mathcal{T}}=N^{-1}$ to target the Bayesian posterior
@@
The matrix divergence $\nabla\cdot\mathrm{M}(z)\in\operatorname{\mathbf{R}}^{d}$ is defined Ma et al. [2015] as
@@
We now cover the most popular SGMCMC algorithms and those implemented in posteriors, all of which are also found in Ma et al. [2015].
@@ B.1 SGLD
The simplest SGMCMC algorithm is stochastic gradient Langevin dynamics [Welling and Teh, 2011], which contains no auxiliary variables (i.e. $z=\theta$) and is represented by $\operatorname{\mathrm{D}}=\mathbb{I}$ and $\operatorname{\mathrm{Q}}=0$. This gives continuous-time SGLD:
@@ B.2 SGHMC
Stochastic gradient Hamiltonian Monte Carlo Chen et al. [2014] extends $z=(\theta,m)$ to include momenta $m$ with Gaussian target distribution
@@ B.3 SGNHT
The stochastic gradient Nosé-Hoover thermostat (SGNHT) algorithm [Ding et al., 2014] uses ideas from thermodynamics [Leimkuhler et al., 2018] to introduce an additional scalar variable $\xi$ that automates the selection of the $\alpha$ friction parameter in SGHMC. We have $z=(\theta,m,\xi)$ targeting  [^2] Note that the SGNHT target distribution for $\xi$ is erroneously claimed as $\frac{1}{2d}(\xi-\alpha)^{2}$ in Ma et al. [2015] and Nemeth and Fearnhead [2021]